Earth absorb 10^22 joule energy every day from the sun what would be the percentage increase in earth mass 6.0×10^24kg due to this
 
 

Grade:12

1 Answers

Arun
5 years ago
 
The sun strikes the Earth with 1022 J of energy every day. Using  E = mc2 
m = 1022/9*1016 = 1.11*105 kg
 
Hope it helps
 
Regards
Arun (askIITians forum expert)